OpenStudy (anonymous):

Teams with varying numbers of players are playing a group card game. The time taken by each team to complete one game is given in the table. What is the relationship between the two variables?

OpenStudy (gorv):

welcome to OS first :)

OpenStudy (gorv):

the Q u posted have insufficient info. Q talk abt the table ..but you didnot posted the table

OpenStudy (anonymous):

umber of Players on the Team Time Taken to Complete Game (in minutes) 1 32 2 26 3 20 4 14 5 8

OpenStudy (gorv):

ok if we take time

OpenStudy (gorv):

32 26 20 14 8

OpenStudy (gorv):

see next term is obtained by subtracting 6 from previous term

OpenStudy (gorv):

let time is y and number of player is x

OpenStudy (gorv):

so y=32+6*(x-1)

OpenStudy (gorv):

x = no of player

OpenStudy (gorv):

if one player is their time= y=32+6*0=32 similarly for other